# On the geometry of spaces of filtrations on local rings

### v1

@@ 3.3.1. Okounkov bodies and continuity along increasing sequences
$\ell(R/\mathfrak{a}_{m})=\sum_{j=1}^{t}(\#\bar{\Gamma}^{(j)}_{m}-\#\Gamma^{(j)}_{m}).$ (3.13)
@@
We proceed to define a valuation $v$ on the quotient field $K:=Q(R)=Q(R^{\prime})$. Fix a regular system of parameters $(x_{1},\ldots,x_{n})$ of $R^{\prime}$ and $\alpha_{1},\ldots,\alpha_{n}\in\mathbb{R}_{>1}$, which are linearly independent over $\mathbb{Q}$. By Cohen’s structure theorem, the $\mathfrak{n}$-completion $(\widehat{R^{\prime}},\hat{\mathfrak{n}})$ of $R^{\prime}$ has a coefficient ring $C$ as a regular local ring. Hence for any $f\in R^{\prime}$ and $d\in\mathbb{Z}_{>0}$, we can write
@@
is independent on the choice of the units and $d$, hence this defines a function $v:K^{\times}\to\mathbb{R}$ by $v(f/g)=v(f)-v(g)$ for any $f,g\in R^{\prime}$. One can check that $v$ is a real valuation. Denote its value group by $\Gamma_{v}:=v(K^{\times})$ and its valuation ring by $(\mathcal{O}_{v},\mathfrak{m}_{v})$ and for any $\lambda\in\mathbb{R}_{\geq 0}$, denote
$\mathfrak{b}_{\lambda}:=\{f\in\mathcal{O}_{v}\mid v(f)\geq\lambda\}\text{ and }\mathfrak{b}_{>\lambda}:=\{f\in\mathcal{O}_{v}\mid v(f)>\lambda\}.$
@@
$\mathfrak{b}_{cm}\cap R\subset\mathfrak{m}^{c_{1}m}\subset\mathfrak{a}_{1}^{m}\subset\mathfrak{a}_{m}.$ (3.14)
@@
We construct $2t$ sequence of semigroups $\Gamma_{,k}^{(j)}$ and $\bar{\Gamma}_{,k}^{(j)}$ associated to $\mathfrak{a}_{\bullet,k}$ as in Proposition 3.13. Let $\Gamma^{(j)}$ and $\bar{\Gamma}^{(j)}$ be the semigroups of $\mathfrak{a}_{\bullet}$. By the proof of Proposition 3.13 it is easy to see that $\bar{\Gamma}_{,k}^{(j)}$ is indeed independent on $k$, that is, $\bar{\Gamma}_{,k}^{(j)}=\bar{\Gamma}^{(j)}$ for any $1\leq j\leq t$ and $k\in\mathbb{Z}_{>0}$. We claim that for any $1\leq j\leq t$,
$\Gamma^{(j)}=\cup_{k}\Gamma_{,k}^{(j)}.$ (3.15)
@@
This show that $(\bm{\beta},m)\in\Gamma_{,K}^{(j)}$. Thus we get $\Gamma^{(j)}\subset\cup_{k}\Gamma_{,k}^{(j)}$ and finishes the proof of the claim.
@@
Let $R=k[\![x]\!]$ with $\mathfrak{m}=(x)$. For $k\in\mathbb{Z}_{>0}$, define $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}$ by
@@
Let $x\in X$ be a klt singularity over a field of characteristic 00. Let $v_{k}\in\mathrm{Val}_{X,x}$, $k\in\mathbb{Z}_{>0}$ be a sequence of valuations such that $A_{X}(v_{k})=1$ and $\widehat{\mathrm{vol}}_{x,X}(v_{k})<V$ for some $V>0$. Possibly passing to a subsequence, there exists an increasing sequence $\{\mathfrak{a}_{\bullet,k}\}\subset\mathrm{Fil}^{s}$ with $\mathfrak{a}_{\lambda,k}\subset\mathfrak{a}_{\lambda}(v_{k})$, which converges to some $v\in\mathrm{Val}_{X,x}$ both weakly and with respect to the $d_{1}$-metric.
@@
For $k,l\in\mathbb{Z}_{>0}$, define
$\mathfrak{a}_{\bullet,k}^{(l)}:=\bigcap_{j=k}^{k+l}\mathfrak{a}_{\bullet}(v_{j}).$
Then for any $k\in\mathbb{Z}_{>0}$, $\{\mathfrak{a}_{\bullet,k}^{(l)}\}_{l\in\mathbb{Z}_{>0}}$ is a decreasing sequence of filtrations. Let
$\mathfrak{a}_{\bullet,k}:=\cap_{l}\mathfrak{a}_{\bullet,k}^{(l)}\subset\mathfrak{a}_{\bullet}(v_{k})$
Clearly the sequence of filtrations $\{\mathfrak{a}_{\bullet,k}\}_{k\in\mathbb{Z}_{>0}}$ is increasing. Let $\chi_{k}=\mathrm{ord}_{\mathfrak{a}_{\bullet,k}}$ be the corresponding norm. By definition, $\chi_{k}=\inf_{j\geq k}v_{j}\leq v$. Since $v_{j}\to v$ weakly, given any $f\in\mathfrak{m}$ and any $\epsilon>0$, there exists $J\in\mathbb{Z}_{>0}$ such that $v_{j}(f)>v(f)-\epsilon$ for any $j\geq J$. Thus
@@
as long as $k\geq J$. This proves $\chi_{k}\to v$ weakly as an increasing sequence. So we may apply Proposition 3.15 to get $\lim_{k\to\infty}d_{1}(\chi_{k},v)=0$, and the proof is finished. ∎
@@ 3.3.2. Normalized volumes and continuity along decreasing sequences
(c.f. [Blu18, Proposition 3.7]) Let $\epsilon,V>0$ and $C\in\mathbb{Z}_{>1}$ be constants. Then for there exists $\Lambda=\Lambda(C,V,\epsilon)\in\mathbb{R}_{>0}$ such that the following holds.
@@
$\mathcal{J}(X;\lambda\cdot\mathfrak{a}_{\bullet})\subset\mathfrak{a}_{\lambda-E},$ (3.16)
@@ 3.4. The toric case and Newton-Okounkov bodies
In this section, we work over a field $\mathbbm{k}$ and follow the notation in [Ful93] for toric varieties. Our goal is to identify the subspace $(\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ of saturated monomial filtrations on a toric singularity with a subspace of the Fréchet-Nikodym-Aronszyan metric on cobounded sets of the dual cone. See [KK14, Section 6] and [Blu18, Section 8] for some relevant constructions.
@@
Recall that there is a $1$-to-$1$ correspondence between toric valuations $v_{u}\in\mathrm{Val}^{\mathrm{toric}}_{R}$ and $u\in\sigma$, given by
@@
Given an $\mathfrak{m}$-primary monomial ideal $\mathfrak{a}$, the Newton polyhedron $P(\mathfrak{a})$ of $\mathfrak{a}$ is the convex hull of $\{m\in\sigma^{\vee}\cap M\mid\chi^{m}\in\mathfrak{a}\}$. Moreover, we have $\mathrm{e}(\mathfrak{a})=\mathrm{vol}(\sigma^{\vee}\backslash P(\mathfrak{a}))$. Recall that the Newton-Okounkov body $P(\mathfrak{a}_{\bullet})$ of a monomial $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{\mathrm{mon}}_{R,\mathfrak{m}}$ is defined to be
@@
which is a convex set such that $\sigma^{\vee}\backslash P(\mathfrak{a}_{\bullet})$ is bounded. It is known that the covolume of the Newton-Okounkov body computes the multiplicity of the filtration.
@@
$\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{vol}(\sigma^{\vee}\backslash P(\mathfrak{a}_{\bullet})).$
@@
For $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{\mathrm{mon}}_{R,\mathfrak{m}}$, $P(\mathfrak{a}_{\bullet})=\cap_{u\in\mathrm{Int}(\sigma)}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
@@
For $m\in\sigma^{\vee}\cap M$ with $\chi^{m}\in\mathfrak{a}_{\lambda}$, we have $\langle u,m\rangle=v_{u}(m)\geq v_{u}(\mathfrak{a}_{\lambda})\geq\lambda v(\mathfrak{a}_{\bullet})$ by definition, that is, $m/\lambda\in H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$. Hence $P(\mathfrak{a}_{\bullet})\subset\cap_{u\in\mathrm{Int}(\sigma)}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
Conversely, assume $p\in\cap_{u\in\mathrm{Int}(\sigma)}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$. Assume $p\notin P(\mathfrak{a}_{\bullet})$. Since $P(\mathfrak{a}_{\bullet})$ is convex, there exists $u\in N_{\mathbb{R}}$ and $b\in\mathbb{R}_{>0}$ such that $\langle u,p\rangle<b$ but $P(\mathfrak{a}_{\bullet})\subset H_{u}(b)$. Since $\sigma^{\vee}\backslash P(\mathfrak{a}_{\bullet})$ is bounded, we know that $u\in\mathrm{Int}(\sigma)$. Now $b>\langle u,p\rangle\geq v_{u}(\mathfrak{a}_{\bullet})$ by the choice of $p$. Hence there exists $\lambda\in\mathbb{R}_{>0}$ and $\chi^{m}\in\mathfrak{a}_{\lambda}$ such that $v_{u}(\chi^{m})=v_{u}(\mathfrak{a}_{\lambda})<\lambda\cdot b$. In particular, $m/\lambda\in P(\mathfrak{a}_{\bullet})\backslash H_{u}(b)$, a contradiction. So $\cap_{u\in\mathrm{Int}(\sigma)}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))\subset P(\mathfrak{a}_{\bullet})$ and the proof is finished. ∎
Next we show that the saturation of a monomial filtration can be computed using only toric valuations.
@@
$\mathfrak{a}_{\bullet}=\cap_{u\in\mathrm{Int}(\sigma)}\mathfrak{a}_{\bullet}(v_{u}/a_{u})$ (3.17)
where $a_{u}:\mathrm{Int}(\sigma)\to\mathbb{R}_{>0}$ is a homogeneous function.
@@
$\mathfrak{a}_{\bullet}\subset\mathfrak{a}^{\prime}_{\bullet}\coloneqq\cap_{u\in\mathrm{Int}(\sigma)}\mathfrak{a}_{\bullet}(\frac{v_{u}}{v_{u}(\mathfrak{a}_{\bullet})})$
by Lemma 2.19. Moreover, $v_{u}(\mathfrak{a}_{\bullet}^{\prime})=v_{u}(\mathfrak{a}_{\bullet})$ since $\mathfrak{a}_{\bullet}\subset\mathfrak{a}^{\prime}_{\bullet}\subset\mathfrak{a}_{\bullet}(v_{u}/v_{u}(\mathfrak{a}_{\bullet}))$. So $P(\mathfrak{a}_{\bullet})=P(\mathfrak{a}^{\prime}_{\bullet})$ by Lemma 3.23, and hence $\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}(\mathfrak{a}^{\prime}_{\bullet})$ by Lemma 3.22. Now $\mathfrak{a}^{\prime}_{\bullet}$ is saturated by Proposition 2.20 again, so by Theorem 2.18, we get $\mathfrak{a}_{\bullet}=\mathfrak{a}^{\prime}_{\bullet}$. ∎
@@
$v_{u}(\cap_{i\in I}\mathfrak{a}_{\bullet,i})=\sup_{i\in I}v_{u}(\mathfrak{a}_{\bullet,i})$ (3.18)
@@
where $+$ denotes the Minkowski sum. Take $p_{i}\in C\cap M_{\mathbb{Q}}$ such that $p_{i}\to p$ and $k_{i}\in\mathbb{Z}_{>0}$ such that $k_{i}p_{i}\in M$. By Lemma 3.23, for any $u\in\mathrm{Int}(\sigma)$, $\langle u,p_{i}\rangle\geq\langle u,p\rangle\geq c^{\prime}_{u}$. Hence $\chi^{k_{i}p_{i}}\in\mathfrak{a}_{k_{i}}\cap\mathfrak{b}_{k_{i}}$by Lemma 3.24, which implies
@@
where $+$ denotes the Minkowski sum. Then there exists $p_{j}\in C\cap M_{\mathbb{Q}}$ such that $p_{j}\to p$. Take $k_{j}\in\mathbb{Z}_{>0}$ such that $k_{j}p_{j}\in M$. Then for any $u\in\mathrm{Int}(\sigma)$, we have $\langle u,p\rangle\leq\langle u,p_{j}\rangle$ for any $u\in\mathrm{Int}(\sigma)$ and
$\langle w,p_{j}\rangle<a_{w}(1-2\delta)+\delta a_{w}=a_{w}(1-\delta).$ (3.19)
@@
Let $\mathcal{P}(\sigma^{\vee})\coloneqq\{P\subsetneq\sigma^{\vee}\mid P\text{ is closed and convex, }\sigma^{\vee}\backslash P\text{ is bounded}\}$. We prove that taking Newton-Okounkov body gives a bijection between $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$ and $\mathcal{P}(\sigma^{\vee})$.
@@
$\mathfrak{a}_{\lambda}(P)\coloneqq\mathrm{span}\{\chi^{m},m\in\sigma^{\vee}\cap M\mid\langle m,u\rangle\geq\lambda\cdot\inf_{p\in P}\langle p,u\rangle\text{ for any }u\in\mathrm{Int}(\sigma)\},$ (3.20)
@@
Thus $\sigma^{\vee}\backslash P(\mathfrak{a}_{\bullet})\subset\cup_{i=1}^{k}\sigma^{\vee}\backslash H_{u_{i}}(1)$, which is bounded. This proves $P:\mathrm{Fil}^{s,\mathrm{mon}}_{X,x}\to\mathcal{P}(\sigma^{\vee})$.
@@
For simplicity, denote $\mathrm{covol}(P)\coloneqq\mathrm{vol}(\sigma^{\vee}\backslash P)$ below. For $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$, we have
@@
$=$ $\mathrm{vol}(P(\mathfrak{a}_{\bullet})\backslash P(\mathfrak{b}_{\bullet}))+\mathrm{vol}(P(\mathfrak{b}_{\bullet})\backslash P(\mathfrak{a}_{\bullet}))$

### v2

@@ 3.3.1. Okounkov bodies and continuity along increasing sequences
We need the following formula, which relates the co-length of a primary ideal to the points of its semigroups.
@@
(3.13) $\ell(R/\mathfrak{a}_{m})=\sum_{j=1}^{t}(\#\bar{\Gamma}^{(j)}_{m}-\#\Gamma^{(j)}_{m})$.
@@
[…]
@@
is independent of the choice of the units and $d$, hence this defines a function $v:K^{\times}\to\mathbb{R}$ by v(f/g)=v(f)v(g)v(f/g)=v(f)-v(g) for any $f,g\in R^{\prime}$. One can check that $v$ is a real valuation. Denote its value group by $\Gamma_{v}:=v(K^{\times})$ and its valuation ring by $(\mathcal{O}_{v},\mathfrak{m}_{v})$ and for any $\lambda\in\mathbb{R}_{\geq 0}$, denote
[…]
@@
(3.14) $\mathfrak{b}_{cm}\cap R\subset\mathfrak{m}^{c_{1}m}\subset\mathfrak{a}_{1}^{m}\subset\mathfrak{a}_{m}.$
@@
It is not hard to see that taking the closed convex cone commutes with taking the union.
@@
Now we are ready to show that the multiplicity is continuous along increasing sequences. The reader may want to compare the following result with similar statements in the global setting, for example, [Dar17, Proposition 6.11] and [BJ25, Lemma 4.7].
@@
We construct $2t$ sequence of semigroups $\Gamma_{,k}^{(j)}$ and $\bar{\Gamma}_{,k}^{(j)}$ associated to $\mathfrak{a}_{\bullet,k}$ as in Proposition 3.13. Let $\Gamma^{(j)}$ and $\bar{\Gamma}^{(j)}$ be the semigroups of $\mathfrak{a}_{\bullet}$. By the proof of Proposition 3.13 it is easy to see that $\bar{\Gamma}_{,k}^{(j)}$ is indeed independent of $k$, that is, $\bar{\Gamma}_{,k}^{(j)}=\bar{\Gamma}^{(j)}$ for any $1\leq j\leq t$ and $k\in\mathbb{Z}_{>0}$. We claim that for any $1\leq j\leq t$,
(3.15) $\Gamma^{(j)}=\cup_{k}\Gamma_{,k}^{(j)}.$
@@
This shows that $(\bm{\beta},m)\in\Gamma_{,K}^{(j)}$. Thus we get $\Gamma^{(j)}\subset\cup_{k}\Gamma_{,k}^{(j)}$ and the proof of the claim is finished.
@@
Let R=k[[x]]R=k[\![x]\!] with $\mathfrak{m}=(x)$. For $k\in\mathbb{Z}_{>0}$, define $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}$ by
@@
Let $x\in X$ be a klt singularity over a field of characteristic 0. Let $v_{k}\in\mathrm{Val}_{X,x}$, $k\in\mathbb{Z}_{>0}$ be a sequence of valuations such that $A_{X}(v_{k})=1$ and $\widehat{\mathrm{vol}}_{x,X}(v_{k})<V$ for some $V>0$. Possibly passing to a subsequence, there exists an increasing sequence $\{\mathfrak{a}_{\bullet,k}\}\subset\mathrm{Fil}^{s}$ with $\mathfrak{a}_{\lambda,k}\subset\mathfrak{a}_{\lambda}(v_{k})$, which converges to some $v\in\mathrm{Val}_{X,x}$ both weakly and with respect to the $d_{1}$-metric.
@@
Let $\chi_{k}\coloneqq\inf_{j\geq k}v_{j}\leq v$. Since $v_{j}\to v$ weakly, given any $f\in\mathfrak{m}$ and any $\epsilon>0$, there exists $J\in\mathbb{Z}_{>0}$ such that $v_{j}(f)>v(f)-\epsilon$ for any $j\geq J$. Thus
@@
as long as $k\geq J$. This proves $\chi_{k}\to v$ weakly as an increasing sequence. So we may apply Proposition 3.15 to get $\lim_{k\to\infty}d_{1}(\chi_{k},v)=0$. By Proposition 2.20, $\mathfrak{a}_{\bullet,k}\coloneqq\mathfrak{a}(\chi_{k})=\cap_{j\geq k}\mathfrak{a}(v_{j})$ is saturated, and the proof is finished. ∎
@@ 3.3.2. Normalized volumes and continuity along decreasing sequences
(c.f. [Blu18, Proposition 3.7]) Let $\epsilon,V>0$ and $C\in\mathbb{Z}_{>1}$ be constants. Then there exists $\Lambda=\Lambda(C,V,\epsilon)\in\mathbb{R}_{>0}$ such that the following holds.
@@
(3.16) $\mathcal{J}(X;\lambda\cdot\mathfrak{a}_{\bullet})\subset\mathfrak{a}_{\lambda-E},$
@@ 3.4. The toric case and Newton-Okounkov bodies
We deal with the toric case in this subsection. Throughout, we work over an algebraically closed field $\mathbbm{k}$ of characteristic 0 and follow the notation in [Ful93] for toric varieties. Our goal is to identify the subspace $(\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ of saturated monomial filtrations on a toric singularity with a subspace of the Fréchet-Nykodym-Aronszajn metric on cobounded sets of the dual cone. See [JM12, Section 8], [KK14, Section 6] and [Blu18, Section 8] for some relevant constructions.
@@
Recall that there is a 11-to-11 correspondence between toric valuations $v_{u}\in\mathrm{Val}^{\mathrm{toric}}_{R}$ and $u\in\sigma$, given by
@@
[…]
@@
[…]
@@
[…]
@@
For $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{\mathrm{mon}}_{R,\mathfrak{m}}$, $P(\mathfrak{a}_{\bullet})=\cap_{u\in\mathrm{Int}(\sigma)\cap N}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
@@
For $m\in\sigma^{\vee}\cap M$ with $\chi^{m}\in\mathfrak{a}_{\lambda}$, we have $\langle u,m\rangle=v_{u}(\chi^{m})\geq v_{u}(\mathfrak{a}_{\lambda})\geq\lambda v(\mathfrak{a}_{\bullet})$ by definition, that is, $m/\lambda\in H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$. Hence $P(\mathfrak{a}_{\bullet})\subset\cap_{u\in\mathrm{Int}(\sigma)\cap N}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
[…]
Next we show that the saturation of a monomial filtration can be computed using only toric divisorial valuations.
@@
(3.17) $\mathfrak{a}_{\bullet}=\cap_{u\in\mathrm{Int}(\sigma)\cap N}\mathfrak{a}_{\bullet}(v_{u}/a_{u})$
where $a_{u}:\mathrm{Int}(\sigma)\to\mathbb{R}_{>0}$ is a homogeneous function for any $u\in\mathrm{Int}(\sigma)\cap N$.
@@
$\mathfrak{a}_{\bullet}\subset\mathfrak{a}^{\prime}_{\bullet}\coloneqq\cap_{u\in\mathrm{Int}(\sigma)\cap N}\mathfrak{a}_{\bullet}(\frac{v_{u}}{v_{u}(\mathfrak{a}_{\bullet})})$
by Lemma 2.19. Moreover, $v_{u}(\mathfrak{a}_{\bullet}^{\prime})=v_{u}(\mathfrak{a}_{\bullet})$ since $\mathfrak{a}_{\bullet}\subset\mathfrak{a}^{\prime}_{\bullet}\subset\mathfrak{a}_{\bullet}(v_{u}/v_{u}(\mathfrak{a}_{\bullet}))$ for any $u\in\mathrm{Int}(\sigma)\cap N$. So $P(\mathfrak{a}_{\bullet})=P(\mathfrak{a}^{\prime}_{\bullet})$ by Lemma 3.23, and hence $\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}(\mathfrak{a}^{\prime}_{\bullet})$ by Lemma 3.22. Now $\mathfrak{a}^{\prime}_{\bullet}$ is saturated by Proposition 2.20 again, and so by Theorem 2.18, we get $\mathfrak{a}_{\bullet}=\mathfrak{a}^{\prime}_{\bullet}$. ∎
@@
(3.18) $v_{u}(\cap_{i\in I}\mathfrak{a}_{\bullet,i})=\sup_{i\in I}v_{u}(\mathfrak{a}_{\bullet,i})$
@@
where ++ denotes the Minkowski sum. Take $p_{i}\in C\cap M_{\mathbb{Q}}$ such that $p_{i}\to p$ and $k_{i}\in\mathbb{Z}_{>0}$ such that $k_{i}p_{i}\in M$. By Lemma 3.23, for any $u\in\mathrm{Int}(\sigma)$, $\langle u,p_{i}\rangle\geq\langle u,p\rangle\geq c^{\prime}_{u}$. Hence $\chi^{k_{i}p_{i}}\in\mathfrak{a}_{k_{i}}\cap\mathfrak{b}_{k_{i}}$by Lemma 3.24, which implies
@@
where ++ denotes the Minkowski sum. Then there exists $p_{j}\in C\cap M_{\mathbb{Q}}$ such that $p_{j}\to p$. Take $k_{j}\in\mathbb{Z}_{>0}$ such that $k_{j}p_{j}\in M$. Then for any $u\in\mathrm{Int}(\sigma)$, we have $\langle u,p\rangle\leq\langle u,p_{j}\rangle$ for any $u\in\mathrm{Int}(\sigma)$ and
(3.19) $\langle w,p_{j}\rangle<a_{w}(1-2\delta)+\delta a_{w}=a_{w}(1-\delta).$
@@
Let $\mathcal{P}(\sigma^{\vee})\coloneqq\{P\subsetneq\sigma^{\vee}\mid P\text{ is closed and convex, }\sigma^{\vee}\backslash P\text{ is bounded}\}$. We prove that taking the Newton-Okounkov body gives a bijection between $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$ and $\mathcal{P}(\sigma^{\vee})$. Note that the following lemma also justifies the statement that “if $P(\mathfrak{a}_{\bullet})=P(\mathfrak{b}_{\bullet})$, then $\mathfrak{a}_{\bullet}$ and $\mathfrak{b}_{\bullet}$ should be regarded as equisingular” in [JM12, Section 8].
@@
(3.20) $\mathfrak{a}_{\lambda}(P)\coloneqq\mathrm{span}\{\chi^{m},m\in\sigma^{\vee}\cap M\mid\langle m,u\rangle\geq\lambda\cdot\inf_{p\in P}\langle p,u\rangle\text{ for any }u\in\mathrm{Int}(\sigma)\},$
@@
[…]
@@
[…]
@@
[…]